# Fairshare Data Pricing via Data Valuation for Large Language Models

### v2

@@ Abstract
Training data is the backbone of large language models (LLMs), yet today’s data markets often operate under exploitative pricing – sourcing data from marginalized groups with little pay or recognition. This paper introduces a theoretical framework for LLM data markets, modeling the strategic interactions between buyers (LLM builders) and sellers (human annotators). We begin with theoretical and empirical analysis showing how exploitative pricing drives high-quality sellers out of the market, degrading data quality and long-term model performance. Then we introduce fairshare, a pricing mechanism grounded in data valuation that quantifies each data’s contribution. It aligns incentives by sustaining seller participation and optimizing utility for both buyers and sellers. Theoretically, we show that fairshare yields mutually optimal outcomes: maximizing long-term buyer utility and seller profit while sustaining market participation. Empirically when training open-source LLMs on complex NLP tasks, including math problems, medical diagnosis, and physical reasoning, fairshare boosts seller earnings and ensures a stable supply of high-quality data, while improving buyers’ performance-per-dollar and long-term welfare. Our findings offer a concrete path toward fair, transparent, and economically sustainable data markets for LLM. Our code will be open sourced.
@@ 1 Introduction
However, the current AI data market operates with limited oversight and is widely criticized for a lack of transparency and fairness in pricing [5, 6, 7, 8]. Data prices are largely low and fail to reflect the quality or effort involved, threatening the sustainability and quality of data supply [9, 10, 11]. In particular, for data sellers (i.e., human annotators or content creators), the prevailing market routinely undervalues their labor, offering compensation that neglects the skill, effort, and downstream value of their contributions. These harms are especially concentrated in low-wage labor markets, where annotators often face overwork, underpayment, and exclusion from decision-making [12]. This reflects a broader ethical concern known as “AI parachuting”, where developers extract data from marginalized communities [13, 14, 15], tieing to wider debates on epistemic injustice and data colonialism [16, 17, 18]. As one civil rights advocate noted on 60 Minutes, “They don’t pay well… they could pay whatever, and have whatever working conditions.” [^1] CBS, 60 Minutes
@@ 2 Related Work
Data Valuation Methods: Data valuation estimates the contribution of training examples to model performance [54]. Examples include influence-function-based methods [55], which estimate data utility via model gradient computations. Variations such of this method enhance efficiency by approximating/bypassing the inverse-Hessian [56], or utilizing lower dimension model gradients (e.g., Datainf [21]), which balance efficiency and accuracy, and make them suitable for the LLM realm. These approaches are effective for LLM training data selection [22, 23] and other applications including toxicity detection [57], memorization analysis [58], and training optimization [59]. Simpler methods like BM25 [30] offer a model-agnostic baseline based on lexical similarity[60, 61].
While exact data value estimation is computationally costly at LLM scale, recent studies show that influence-based approximations correlate meaningfully with actual performance outcomes [62, 63]. For example, Jiao et al. [62] shows that influence-based methods have high correlation between their estimated data valuation scores and the oracle value, which is obtained through model re-training.
@@ 3.1 Data Market Definition
We formalize the LLM data market by modeling the decision-making processes of sellers and buyers in a Stackelberg game [35, 45, 48]. In this setup, sellers first set prices by anticipating buyer demand. Buyers then respond by selecting datasets to maximize their utility. [^2] Utility is defined as a general economic concept that measures an individual’s welfare, benefit, or satisfaction [24, 25], for example, monetary value, model fairness, or other domain-specific benefits. Each player makes decisions based solely on the current state of the market, without considering future updates or interactions.
@@
The value $v_{k}$ can be estimated using a data valuation method (e.g., Influence Function [55], DataInf [21], etc.). Once the data valuation score $v_{k}$ is estimated, the corresponding utility gain $u_{k}$ can be derived as a function of $v_{k}$. We use a task-specific mapping between data value $v_{k}$ and utility $u_{k}:v_{k}\rightarrow\mathbb{R}_{\geq 0}$. Appendix C lists common mappings between $u_{k}$ and $v_{k}$ in downstream tasks.
@@
This utility gain reflects the downstream value, resulting from performance improvements in the buyer’s model after training on the dataset bundle $\boldsymbol{x}$ [^3] Datasets may exhibit dependencies. The set utility $u_{k}(\boldsymbol{x})$ is not necessarily additive over individuals. The net utility is defined as:
@@
Seller’s Decision-Making. When the seller $S_{j}$ offers dataset $D_{j}$, it sets a price to maximize its net profit, defined as (i) the anticipated sales from all buyers, minus (ii) a fixed cost to create the dataset. Anticipated sales from buyers are estimated by solving the previous buyer’s purchasing problem (Equation 2) known in the full transparent market. Formally, the seller’s net profit function is:
@@
where $\tilde{\boldsymbol{x}}^{k,N}_{j}(\boldsymbol{p})$ is $j$-th entry of buyer $B_{k}$’s decision vector, indicating if $B_{k}$ purchases $D_{j}$ ($j$-th entry is $1$) or not ($j$-th entry is 00), given $\boldsymbol{p}$ the prices of all datasets in the market; $c_{j}$ denotes the cost. The seller $S_{j}$ solves the following pricing problem:
@@
where $p^{*}_{j}$ is the optimal price and $r(p_{j})\geq 0$ ensures that seller $S_{j}$’s net profit must be non-negative.
In the above formulation, buyer purchase decisions depend on prices and data value; seller pricing decisions depend on anticipated sales from buyers. Together, these decisions define the equilibrium dynamics of a transparent and incentive-aligned LLM data market. This formulation can be extended to a royalty-based scheme as presented in Appendix B.
@@ 3.2 Exploitative Pricing: A Lose-Lose Outcome
Assumption 1 models how the seller respond to exploitative pricing over time. It states that when offered a price below its ideal level, the seller become less likely to stay in the market, consistent with prior studies [65, 66]. For any exploitative pricing, the participation probability $P_{T}$ declines multiplicatively, leading to a sustained and irreversible drop in engagement.
@@
A higher discount factor $\delta$ indicates greater emphasis on future utility. Assumption 2 places a lower bound on $\delta$, ensuring that the buyer sufficiently values future gains when making decisions.
Assumption 2 assumes the buyer values both immediate and future utility gains from acquiring training data. Major LLM developers invest in large-scale data acquisition and model training in expectation of long-term gains in performance, deployment value, and commercial returns [67, 68, 69, 70].
@@
With Assumptions 1, 1.1 and 2, any exploitative pricing (i.e., $p_{t}<p_{t}^{*},\forall t$) will only maximize cumulative utility within a finite horizon – after which it is strictly suboptimal.
Lemma 1 reveals a fundamental limitation: any exploitative pricing strategy is only optimal for a finite time. Over time, it becomes strictly suboptimal, due to the declining seller participation. Thus, no exploitative strategy can maximize cumulative utility in the long run. (See Appendix D for the proof, shows that exploitative pricing yields a suboptimal value function.)
@@
In addition to our theoretical findings, we run a simplified dynamic market simulation with one buyer and one seller, making sequential decisions. The utility of the dataset and buyer’s budget varies randomly over time, and seller participation follows $\pi(p_{t},p_{t}^{*})=p_{t}/p_{t}^{*}$. Figures 2(a) and 2(b) provides consistent evidence that exploitative pricing causes rapid seller exit and an immediate shortage of training data. This dynamic closely resembles the classic “market for lemons” problem, where underpricing drives out high-quality supply, ultimately leading to market collapse [8].
@@ 3.3 Fairshare Pricing: A Win-Win for Sellers and Buyers
Seller $S_{j}$’s optimal price for $D_{j}$ is characterized as one of buyers’ MWP (see Appendix D for proof):
@@
As shown in Figures 2(c) and 2(d), we simulate a data market with $2$ sellers and $2$ buyers, where players make one-shot decisions. Seller $S_{1}$ sets its price first, followed by seller $S_{2}$. Each plot shows each seller’s profit function $r(p_{j})$, with breakpoints at each buyer’s MWP. Pricing above a buyer’s MWP leads to a drop in sales. Thus, the seller’s optimal price aligns with the MWP that yields maximum revenue.
Buyer-Side: Fairshare Is Overall Optimal. We now show that the fairshare price $p_{j}^{*}$, derived from seller optimization, is also optimal from the buyer’s perspective. As established in Section 3.2, exploitative pricing leads to persistent seller exit. In contrast, fairshare pricing ensures full participation and maximizing the buyer’s overall utility. To formalize this, we consider the same infinite-horizon setting in Section 3.2 with a single buyer and seller. Notably, the fairshare price is the ideal price for sellers as it maximizes its profit. Therefore, under fairshare pricing, the sellers will maintain sustained participation. Then, we obtain:
@@
The fairshare price for the seller $S$ under LLM data framework is $p^{*}_{t}=\min\{u_{t},b_{t}\},\forall t$. With Assumptions 1, 1.1 and 2, $p^{*}_{t}$ is the optimal pricing strategy that maximizes buyer’s expected cumulative utility (Equation 8).
@@
In Figure 3, we develop a synthetic simulation illustrating that a buyer’s optimal pricing strategy rapidly converges to the fairshare price. Consider a simulation with players receive constant utility $u_{t}=2$ and budget $b_{t}=1,\forall t$, with the discount factor $\delta=0.95$, so the seller’s optimal price remains $p^{*}_{t}=1$. At each time step, the buyer selects a price that maximizes cumulative net utility. The buyer begins with a low, exploitative price at $T=1$, but by $T=5$, converges to the fairshare price and maintains it thereafter. This illustrates our core theoretical insight: fair pricing emerges as the optimal long-run strategy when buyers account for overall market sustainability.
@@ 4.1 Experimental Setup
In our experiments, a buyer is equipped with a single LLM, and each seller owns one data sample. Buyers seek to buy training data to improves task-specific model performance (e.g.,math problem solving, medical diagnosis, or physical reasoning), which in turn increases their utility. (We assume an affine mapping between performance and utility; see Section C.1.)
@@
Sellers and Datasets. We focus on challenging, human-annotated tasks: MathQA and GSM8K [27, 75] for math, MedQA [28] for medical diagnosis, and PIQA [29] for physical reasoning [76, 77, 78]. Table 1 in Appendix F shows dataset splits and examples. We use the training splits as seller data and simulate the market dynamics from Section 3, treating each task as a market scenario.
@@ 4.2 LLM Data Market Experiments
Figure 4: (1) buyer’s cumulative utilities with high- (Figure 4(a)) and low-budget buyer (Figure 4(b)), and (2) sellers’ average cumulative profits (Figure 4(c)) and active seller numbers (Figure 4(d)) over $100$ time periods. Pythia-1b; MedQA; Groups: (1) fairshare, (2) reduced, (3) random, and (4) exploitative.
@@
Market Design: Following the general setups in Section 4.1, we simulate a market with $2$ buyers and $10$ sellers over multiple time steps. To examine buyers with varying resources, we include a high-budget buyer (well-funded LLM builder) and a low-budget buyer (under-resourced one). Each buyer’s budget is randomly drawn from distributions with different mean. At each time step, (1) sellers arrive sequentially with a new dataset ( $300$ data samples) at a fixed price, and then (2) once all sellers arrive, buyers make purchases by solving Equation 2. Full details are in Section E.2.
Participation Function: Following Assumptions 1, 1.1 and 2, we simulate $100$ time steps with a discount factor $\delta=0.98$. We also run experiments with different value of $\delta$ showing consistent and robust results. (See Appendix E). Seller participation probability is defined as $\pi(p_{j,t},p^{*}_{j,t})=p_{j,t}/p^{*}_{j,t}$ for its simplicity and compliance with Assumptions 1 and 1.1. Sellers receiving exploitative pricing ( $p_{j,t}<p^{*}_{j,t}$) are less likely to participate.
@@
Figure 4 compares the overall welfare outcomes of different pricing methods for buyers and sellers using Pythia-1b on the MedQA task. Results on MathQA and PIQA (with Pythia-410m and Llama-3.2-Instruct-1b) in Appendix F show similar patterns.
@@
Figure 5: Purchased datasets for the buyer with high budget (Figure 5(a)) and low budget (Figure 5(b)) over $100$ time periods. Pythia-1b; MedQA.
Exploitative Pricing Leads to Lose-Lose Outcomes: The exploitative pricing method, which sets uniformly low, fixed prices to reflect real-world practices ([11]), offers LLM developers short-term utility gains (see Figures 4(a) and 4(b)). However, this approach systematically undervalues datasets and unfairly compensates annotators. As a result, sellers exit the market over time, triggering a collapse in data supply. Even well-funded LLM developers are unable to source sufficient data – hindering the advancement of LLMs.
Fairshare Pricing Leads to Win-Win Outcomes: For sellers, fairshare pricing consistently yields the highest profit over time (Figure 4(c)), aligning with predictions from our theoretical model (Section 3.1). For buyers, fairshare pricing framework is particularly effective for the high-budget buyer (Figure 4(a)), maximizing long-term utility. However, the low-budget buyer (Figure 4(b)) experiences reduced short-term utility in exchange for long-term gains. Its limited budget prevents them from fully leveraging the increased dataset supply ensured by fairshare pricing, making other low-pricing methods initially more appealing. Yet, in the long run, fairshare pricing sustains seller participation, ensuring data supply.

### v3

@@ Abstract
Training data is the backbone of large language models (LLMs), yet today’s data markets often operate under exploitative pricing – sourcing data from marginalized groups with little pay or recognition. This paper introduces a theoretical framework for LLM data markets, modeling the strategic interactions between buyers (LLM builders) and sellers (human annotators). We begin with theoretical and empirical analysis showing how exploitative pricing drives high-quality sellers out of the market, degrading data quality and long-term model performance. Then we introduce fairshare, a pricing mechanism grounded in data valuation that quantifies each data’s contribution. It aligns incentives by sustaining seller participation and optimizing utility for both buyers and sellers. Theoretically, we show that fairshare yields mutually optimal outcomes: maximizing long-term buyer utility and seller profit while sustaining market participation. Empirically when training open-source LLMs on complex NLP tasks, including math problems, medical diagnosis, and physical reasoning, fairshare boosts seller earnings and ensures a stable supply of high-quality data, while improving buyers’ performance-per-dollar and long-term welfare. Our findings offer a concrete path toward fair, transparent, and economically sustainable data markets for LLM.
@@ 1 Introduction
However, the current AI data market operates with limited oversight and is widely criticized for a lack of transparency and fairness in pricing [5, 6, 7, 8]. Data prices are largely low and fail to reflect the quality or effort involved, threatening the sustainability and quality of data supply [9, 10, 11]. In particular, for data sellers, such as human annotators or content creators, the prevailing market routinely undervalues their labor, offering compensation that neglects the skill, effort, and downstream value of their contributions. These harms are especially concentrated in low-wage labor markets, where annotators often face overwork, underpayment, and exclusion from decision-making [12]. This reflects a broader ethical concern known as “AI parachuting”, where developers extract data from marginalized communities [13, 14, 15], tieing to wider debates on epistemic injustice and data colonialism [16, 17, 18]. As one civil rights advocate noted on 60 Minutes, “They don’t pay well… they could pay whatever, and have whatever working conditions.” [^1] CBS, 60 Minutes
@@ 2 Related Work
Data Valuation Methods: Data valuation estimates the contribution of training examples to model performance [54]. Examples include influence-function-based methods [55], which estimate data utility via model gradient computations. Variations such of this method enhance efficiency by approximating/bypassing the inverse-Hessian [56], or utilizing lower dimension model gradients (e.g., Datainf [21]), which balance efficiency and accuracy, and make them suitable for the LLM realm. Beyond the influence of loss functions, previous research [57] has examined data valuation with respect to fractional utility functions.
Data valuation approaches are effective for LLM training data selection [22, 23, 58] and other applications including toxicity detection [59], memorization analysis [60], training optimization [61], and and active sampling [62]. Shapley-based method is another direction [63, 64, 65, 36], such as CS-Shaply[66], In-run Shapley[67], DU-Shapley[68], etc. Simpler methods like BM25 [30] offer a model-agnostic baseline based on lexical similarity[69, 70]. While exact data value estimation is computationally costly at LLM scale, recent studies show that influence-based approximations correlate meaningfully with actual performance outcomes [71, 72]. For example, Jiao et al. [71] shows that influence-based methods have high correlation between their estimated data valuation scores and the oracle value, which is obtained through model re-training.
@@ 3.1 Data Market Definition
We formalize the LLM data market by modeling the decision-making processes of sellers and buyers in a non-cooperative Stackelberg game [35, 45, 48]. In this setup, sellers first set prices by anticipating buyer demand. Buyers then respond by selecting datasets to maximize their utility. [^2] Utility is defined as a general economic concept that measures an individual’s welfare, benefit, or satisfaction [24, 25], for example, monetary value, model fairness, or other domain-specific benefits. Each player makes decisions based solely on the current state of the market, without considering future updates or interactions.
Notably, our choice of a non-cooperative Stackelberg game fits the LLM data market’s reality: many independent, self-interested transactions where binding grand coalitions are infeasible, such as gig-work platforms [18]. In contrast, cooperative bargaining targets few-party, contractible surplus splits, which is a different regime and question. In this non-cooperative game, we model fairness behaviorally via seller participation/exit that keeps the model tractable, and the leader–follower timing mirrors real price setting and demand response.
@@
The value $v_{k}$ can be estimated using a data valuation method (e.g., Influence Function [55], DataInf [21], etc.). Once the data valuation score $v_{k}$ is estimated, the corresponding utility gain $u_{k}$ can be derived as a function of $v_{k}$. We use a task-specific mapping between data value $v_{k}$ and utility $u_{k}:v_{k}\rightarrow\mathbb{R}_{\geq 0}$. Appendix˜C lists common mappings between $u_{k}$ and $v_{k}$ in downstream tasks.
@@
This utility gain reflects the downstream value, resulting from performance improvements in the buyer’s model after training on the dataset bundle $\boldsymbol{x}$. [^3] Datasets may exhibit dependencies. The set utility $u_{k}(\boldsymbol{x})$ is not necessarily additive over individuals. The net utility is defined as:
@@
Seller’s Decision-Making. When the seller $S_{j}$ offers dataset $D_{j}$, it sets a price to maximize its net profit, defined as (i) the anticipated sales from all buyers, minus (ii) a fixed cost to create the dataset. Anticipated sales from buyers are estimated by solving the previous buyer’s purchasing problem (Equation˜2) known in the full transparent market. Formally, the seller’s net profit function is:
@@
where $\tilde{\boldsymbol{x}}^{k,N}_{j}(\boldsymbol{p})$ is $j$-th entry of buyer $B_{k}$’s decision vector, indicating if $B_{k}$ purchases $D_{j}$ ($j$-th entry is $1$) or not ($j$-th entry is 0), given $\boldsymbol{p}$ the prices of all datasets in the market; $c_{j}$ denotes the cost. The seller $S_{j}$ solves the following pricing problem:
@@
where $p^{*}_{j}$ is the optimal price and $r(p_{j})\geq 0$ ensures that seller $S_{j}$’s net profit must be non-negative. It is noted that we assume that selling data for annotator is profitable, i.e., $p^{*}_{j}\geq c_{j}$.
In the above formulation, buyer purchase decisions depend on prices and data value; seller pricing decisions depend on anticipated sales from buyers. Together, these decisions define the equilibrium dynamics of a transparent and incentive-aligned LLM data market. This formulation can be extended to a royalty-based scheme as presented in Appendix˜B.
@@ 3.2 Exploitative Pricing: A Lose-Lose Outcome
Assumption˜1 models how the seller respond to exploitative pricing over time. It states that when offered a price below its ideal level, the seller become less likely to stay in the market, consistent with prior studies [74, 75]. For any exploitative pricing, the participation probability $P_{T}$ declines multiplicatively, leading to a sustained and irreversible drop in engagement.
@@
A higher discount factor $\delta$ indicates greater emphasis on future utility. Assumption˜2 places a lower bound on $\delta$, ensuring that the buyer sufficiently values future gains when making decisions.
Assumption˜2 assumes the buyer values both immediate and future utility gains from acquiring training data. Major LLM developers invest in large-scale data acquisition and model training in expectation of long-term gains in performance, deployment value, and commercial returns [76, 77, 78, 79].
@@
With Assumptions˜1, 1.1 and 2, any exploitative pricing (i.e., $p_{t}<p_{t}^{*},\forall t$) will only maximize cumulative utility within a finite horizon – after which it is strictly suboptimal.
Lemma˜1 reveals a fundamental limitation: any exploitative pricing strategy is only optimal for a finite time. Over time, it becomes strictly suboptimal, due to the declining seller participation. Thus, no exploitative strategy can maximize cumulative utility in the long run. (See Appendix˜D for the proof, shows that exploitative pricing yields a suboptimal value function.)
@@
In addition to our theoretical findings, we run a simplified dynamic market simulation with one buyer and one seller, making sequential decisions. The utility of the dataset and buyer’s budget varies randomly over time, and seller participation follows $\pi(p_{t},p_{t}^{*})=p_{t}/p_{t}^{*}$. Figures˜2(a) and 2(b) provides consistent evidence that exploitative pricing causes rapid seller exit and an immediate shortage of training data. This dynamic closely resembles the classic “market for lemons” problem, where underpricing drives out high-quality supply, ultimately leading to market collapse [8].
@@ 3.3 Fairshare Pricing: A Win-Win for Sellers and Buyers
Seller $S_{j}$’s optimal price for $D_{j}$ is characterized as one of buyers’ MWP (see Appendix˜D for proof):
@@
As shown in Figures˜2(c) and 2(d), we simulate a data market with $2$ sellers and $2$ buyers, where players make one-shot decisions. Seller $S_{1}$ sets its price first, followed by seller $S_{2}$. Each plot shows each seller’s profit function $r(p_{j})$, with breakpoints at each buyer’s MWP. Pricing above a buyer’s MWP leads to a drop in sales. Thus, the seller’s optimal price aligns with the MWP that yields maximum revenue.
Buyer-Side: Fairshare Is Overall Optimal. We now show that the fairshare price $p_{j}^{*}$, derived from seller optimization, is also optimal from the buyer’s perspective. As established in Section˜3.2, exploitative pricing leads to persistent seller exit. In contrast, fairshare pricing ensures full participation and maximizing the buyer’s overall utility. To formalize this, we consider the same infinite-horizon setting in Section˜3.2 with a single buyer and seller. Notably, the fairshare price is the ideal price for sellers as it maximizes its profit. Therefore, under fairshare pricing, the sellers will maintain sustained participation. Then, we obtain:
@@
The fairshare price for the seller $S$ under LLM data framework is $p^{*}_{t}=\min\{u_{t},b_{t}\},\forall t$. With Assumptions˜1, 1.1 and 2, $p^{*}_{t}$ is the optimal pricing strategy that maximizes buyer’s expected cumulative utility (Equation˜8).
@@
In Figure˜3, we develop a synthetic simulation illustrating that a buyer’s optimal pricing strategy rapidly converges to the fairshare price. Consider a simulation with players receive constant utility $u_{t}=2$ and budget $b_{t}=1,\forall t$, with the discount factor $\delta=0.95$, so the seller’s optimal price remains $p^{*}_{t}=1$. At each time step, the buyer selects a price that maximizes cumulative net utility. The buyer begins with a low, exploitative price at $T=1$, but by $T=5$, converges to the fairshare price and maintains it thereafter. This illustrates our core theoretical insight: fair pricing emerges as the optimal long-run strategy when buyers account for overall market sustainability.
Next, we also explore the role of the discount factor $\delta$ in our framework:
Lemma 4 (The Trade-Off Threshold Is Increasing as $\delta$ Decreases).
And $t^{*}$ is increasing as $\delta$ increases.
We run detailed robustness check with different values of $\delta$ for our experiments in Section˜E.2.
@@ 4.1 Experimental Setup
In our experiments, a buyer is equipped with a single LLM, and each seller owns one data sample. Buyers seek to buy training data to improves task-specific model performance (e.g.,math problem solving, medical diagnosis, or physical reasoning), which in turn increases their utility. (We assume an affine mapping between performance and utility; see Section˜C.1.)
@@
Sellers and Datasets. We focus on challenging, human-annotated tasks: MathQA and GSM8K [27, 84] for math, MedQA [28] for medical diagnosis, and PIQA [29] for physical reasoning [85, 86, 87]. Table˜1 in Appendix˜F shows dataset splits and examples. We use the training splits as seller data and simulate the market dynamics from Section˜3, treating each task as a market scenario.
@@ 4.2 LLM Data Market Experiments
Figure 4: (1) buyer’s cumulative utilities with high- (Figure˜4(a)) and low-budget buyer (Figure˜4(b)), and (2) sellers’ average cumulative profits (Figure˜4(c)) and active seller numbers (Figure˜4(d)) over $100$ time periods. Pythia-1b; MedQA; Groups: (1) fairshare, (2) reduced, (3) random, and (4) exploitative.
@@
Market Design: Following the general setups in Section˜4.1, we simulate a market with $2$ buyers and $10$ sellers over multiple time steps. To examine buyers with varying resources, we include a high-budget buyer (well-funded LLM builder) and a low-budget buyer (under-resourced one). Each buyer’s budget is randomly drawn from distributions with different mean. At each time step, (1) sellers arrive sequentially with a new dataset ( $300$ data samples) at a fixed price, and then (2) once all sellers arrive, buyers make purchases by solving Equation˜2. Full details are in Section˜E.2.
Participation Function: Following Assumptions˜1, 1.1 and 2, we simulate $100$ time steps with a discount factor $\delta=0.98$. We also run experiments with different value of $\delta$ showing consistent and robust results. (See Appendix˜E). Seller participation probability is defined as $\pi(p_{j,t},p^{*}_{j,t})=p_{j,t}/p^{*}_{j,t}$ for its simplicity and compliance with Assumptions˜1 and 1.1. Sellers receiving exploitative pricing ( $p_{j,t}<p^{*}_{j,t}$) are less likely to participate.
@@
Figure˜4 compares the overall welfare outcomes of different pricing methods for buyers and sellers using Pythia-1b on the MedQA task. Results on MathQA and PIQA (with Pythia-410m and Llama-3.2-Instruct-1b) in Appendix˜F show similar patterns.
@@
Figure 5: Purchased datasets for the buyer with high budget (Figure˜5(a)) and low budget (Figure˜5(b)) over $100$ time periods. Pythia-1b; MedQA.
Exploitative Pricing Leads to Lose-Lose Outcomes: The exploitative pricing method, which sets uniformly low, fixed prices to reflect real-world practices ([11]), offers LLM developers short-term utility gains (see Figures˜4(a) and 4(b)). However, this approach systematically undervalues datasets and unfairly compensates annotators. As a result, sellers exit the market over time, triggering a collapse in data supply. Even well-funded LLM developers are unable to source sufficient data – hindering the advancement of LLMs.
Fairshare Pricing Leads to Win-Win Outcomes: For sellers, fairshare pricing consistently yields the highest profit over time (Figure˜4(c)), aligning with predictions from our theoretical model (Section˜3.1). For buyers, fairshare pricing framework is particularly effective for the high-budget buyer (Figure˜4(a)), maximizing long-term utility. However, the low-budget buyer (Figure˜4(b)) experiences reduced short-term utility in exchange for long-term gains. Its limited budget prevents them from fully leveraging the increased dataset supply ensured by fairshare pricing, making other low-pricing methods initially more appealing. Yet, in the long run, fairshare pricing sustains seller participation, ensuring data supply.